OpenStudy (anonymous):

Express in simplified form the value of integral of (2e^x + 4)dx from 0 to ln(2)

OpenStudy (mathmale):

\[\int\limits_{0}^{\ln 2}(2e ^{x}+4)dx=2\int\limits_{0}^{\ln 2}e ^{x}dx +4\int\limits_{0}^{\ln 2}dx\] Can you, K.O, take the integration from this point to completion?